If the stone in fig A enters the water at 41.6 ms-1, and takes 0.6s to travel the 3 meters to the bottom of the pond, what is its average acceleration in the pond water?


1
Expert's answer
2021-01-05T12:04:34-0500

The distance travelled under a constant acceleration is given as follows:



"d = v_0t -\dfrac{at^2}{2}"

where "d = 3m" is the depth of the pond, "t = 0.6s" is time of movement, "v_0 = 41.6m\/s" is the initial velocity, and "a" is the required average acceleration. Expressing "a", find:



"a = \dfrac{2v_0}{t} - \dfrac{2d}{t^2}\\
a = \dfrac{2\cdot 41.6}{0.6} - \dfrac{2\cdot 3}{0.6^2} = 122m/s^2"

Answer. 122 m/s^2.